What is the answer to 4b-1/7=5/7?
Answer by candicem(3)

1.Add 1/7 to both side to cancel the -1/7.
you get 4b=6/7
2.Divide both sides by 4 to get b by itself.
4b/4=(6/7)/4
b=(6/7)/4
3.Solve
b=3/14